Express in kg: 200 g Express in kg: 200 g Express in kg: 200 g

Express in kg: 
200 g 
Express in kg: 
200 g 
Express in kg: 
200 g 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that,
= 1 kg = 1000 g
Then,
= 1 g = (1/1000) kg
= 200 g
= (200/1000) kg
= (2/10)
= 0.2 kg

Thanks